OpenStudy (anonymous):

A referee tosses a coin straight up with an initial velocity of 19 ft per second from a height of 6 ft. What is the maximum height of the coin? if he doesnt catch it when will it hit the ground?

OpenStudy (anonymous):

@AnswerMyQuestions @Data_LG2 @mathmate @math&ing001

OpenStudy (mathmate):

Maximum height is given by \(h=\large \frac{v^2}{2g}\). You can get the formula by equation potential and kinetic energies. A free fall object starting from rest follows the rules: Going up: v=u+at (v=final velocity, u=initial velocity) Here u=19, v=0, a=-9.8 Solve for t. Going down, S=ut+(1/2)at^2 (S=distance travelled) Here, u=0 (falling from rest) a=-9.8 S=-(h+6) (use h calculated earlier) Solve for t. Add together the two t's to get total time.